OpenStudy (anonymous):

please show your work. F(x)=2-3x^2 and g(x)=x+2. find f 0 g and g 0 f

OpenStudy (anonymous):

fog = 2-3(x+2)^2

OpenStudy (anonymous):

gof = (2-3x^2) +2

OpenStudy (anonymous):

you can simplify urself

OpenStudy (anonymous):

how do I do it?

OpenStudy (anonymous):

caqn you walk me through the steps

OpenStudy (anonymous):

fog = 2 -3(x+2)(x+2) etc gof = 2 -3x^2 +2 = -3x^2

OpenStudy (anonymous):

if fo g the g function takes the place of the x as it is f(g(x)) in the gof function we place the f function for the x in g

OpenStudy (anonymous):

ok